without graphing find the slope and the vertical intercept of the line modeled by each equation
y = 1/3 - 2/3 x
y = 7x
y = - 1/8x + 7/8

Answer by sprolden(40)

without graphing find the slope and the vertical intercept of the line modeled by each equation
y = 1/3 - 2/3 x
y = 7x
y = - 1/8x + 7/8
1- The equation of a line is y = mx+b, where m=slope (also the coefficient of x) and b=y-intercept.
2- In the first equation, notice that the slope would be -2/3 since it is the coefficient of x. The y-intercept would be 1/3.
3- In the second equation, the slope is 7 and the y-intercept is 0.
4- In the third equation, the slope is -1/8 and the y-intercept is 7/8.